OpenStudy (anonymous):

At 2:00, Braden rented a mountain bike from the Slick Rock Company. He returned the bike at 5:00, after cycling 18 mi. Braden paid $12 for the rental. a) Find Braden"s average speed, in miles per hour. b) Find the rental rate, in dollars per hour. c) Find the rental rate, in dollars per mile.

OpenStudy (anonymous):

a) 2:00 to 5:00 is 3 hours. If Branden cycled 18 miles in 3 hours his average speed is 6mph as speed=distance over time (18/3=6). b) Branden cycled for 3 hours and paid $12, therefore the rental rate per hour = $12/3 = $4. c) Rental rate per mile = $12/18 = $0.666666....

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."